OpenStudy (risingtotop):

Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. (5 points) Part B: How many liters of apple juice are present in the jar? Show your work. (5 points)

OpenStudy (risingtotop):

@3mar

OpenStudy (3mar):

Well, I am here.

OpenStudy (risingtotop):

Hello

OpenStudy (3mar):

Hello!

OpenStudy (3mar):

3 minutes please

OpenStudy (risingtotop):

alright

OpenStudy (3mar):

thanks

OpenStudy (3mar):

Well, I am here.

OpenStudy (3mar):

Medal before we begin!! Nice stimulation!

OpenStudy (risingtotop):

i think part A equation is 0.75/1=x

OpenStudy (3mar):

What if I think other than this? ;) hahah

OpenStudy (risingtotop):

is my equation right?

OpenStudy (3mar):

Unfortunately, no Write an equation in one variable that can be used to find """"the total number of liters""">>1 lit of apple juice and water in the jar>>>>>0.75+0.25 how did you define the variable???

OpenStudy (3mar):

what is the definition of X???

OpenStudy (3mar):

@RisingToTop Do you follow?

OpenStudy (risingtotop):

@3mar

OpenStudy (3mar):

Where are you?

OpenStudy (risingtotop):

Here I was doing something

OpenStudy (risingtotop):

sorry

OpenStudy (risingtotop):

and yes i follow

OpenStudy (risingtotop):

0.75+0.25x

OpenStudy (3mar):

what is the definition of X??? can you define x for me, then type in your equation?

OpenStudy (risingtotop):

x is "rest of water"

OpenStudy (3mar):

either you write x, or write 0.25lit, because they are the same.

OpenStudy (risingtotop):

0.25

OpenStudy (3mar):

so where would you type x?

OpenStudy (risingtotop):

0.25x

OpenStudy (3mar):

what I mean is : when you name x is the "amount of rest of water" so the equation would be: 0.75+x=1

OpenStudy (risingtotop):

ooooo

OpenStudy (3mar):

Did you get it?

OpenStudy (risingtotop):

yes i did now

OpenStudy (risingtotop):

Actually i understand perfectly

OpenStudy (3mar):

if the contrast: x represents the amount of apple juice, it will be x+0.25=1 while in both cases the right side represents the total amount of the juice in the jar

OpenStudy (3mar):

Praise for Allah. I am so happy for hearing that!

OpenStudy (risingtotop):

Part A: o.75+x=1 Part B: 0.25+5=1 @3mar

OpenStudy (3mar):

Is that logic? part B? is 5.25=1 ???!

OpenStudy (risingtotop):

o.25+x

OpenStudy (3mar):

He wants what? ""How many liters of apple juice are present in the jar?"" i.e separate the ""apple juice" amount in one side........toooooo easyyyyy!

OpenStudy (risingtotop):

1.75

OpenStudy (3mar):

This is Part A: o.75+x=1 isolate x. Done!

OpenStudy (3mar):

Bring x in one side alone and the other terms in the other side!

OpenStudy (risingtotop):

.75 sorry!

OpenStudy (3mar):

He wants the form: \[0.75=1-x\]

OpenStudy (risingtotop):

i dont get it

OpenStudy (risingtotop):

so whats part a and part B??

OpenStudy (3mar):

part a is such as you got...correct " find the total number of liters of apple juice and water" part b is "How many liters of apple juice": which is 0.75=total amount-water contert 0.75=1-x Is not it?

OpenStudy (risingtotop):

1.75 L of apple juice are present??!!

OpenStudy (3mar):

yes

OpenStudy (3mar):

but you should show x in the equation

OpenStudy (risingtotop):

We are done with this problem yes?

OpenStudy (3mar):

YYYYYYYYYYEEEEEEEEEEEEESSSSSSSSSSSSSS!

OpenStudy (3mar):

Congratulations!

OpenStudy (3mar):

Thank you for learning!! Thanks for your patience!

OpenStudy (3mar):

The most important thing is : did you got what you are looking for?

OpenStudy (risingtotop):

yes

OpenStudy (3mar):

Thanks for Allah.

OpenStudy (3mar):

Any more questions?